OpenStudy (anonymous):

Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form.y – 4 = 2(x + 3) y + 4 = 2(x – 3) y + 3 = 2(x – 4) y – 3 = 2(x + 4)

OpenStudy (jdoe0001):

so let's use the point-slope form to get it => \(\huge y-y_1=m(x-x_1)\) where "m" is the slope there, plug your values in, what do you get?

OpenStudy (jdoe0001):

you are given 2 values, so \(\large (-3, 4) \implies (x_1, \ y_1)\)

OpenStudy (texaschic101):

do you understand what jdoe is saying ?

OpenStudy (texaschic101):

Let me get you started... y - y1 = m(x - x1) using slope(m)2 and points(-3,4) y - 4 = 2(x -(-3) Now don't forget to distribute,and remember two negatives multiplied together make a positive.